# Algorithms and complexity for monitoring edge-geodetic sets in graphs

### v1

@@ 1 Introduction
In the introductory paper on MEG-sets [20] (whose extended version is [15]), the parameter $\operatorname{meg}(G)$ was determined for various graph families such as trees, cycles, unicyclic graphs, complete graphs, complete multipartite graphs, hypercubes, grids. This parameter was further studied for Mycielskian graph classes [30], line graphs [3], Cartesian and strong products of graphs [24], corona products [15], cluster, lexicographic product, direct product and the join operation [35]. The effects of clique sum and subdivision on $\operatorname{meg}(G)$ are also explored in [18]. The relation between $\operatorname{meg}(G)$ and other parameters of graphs has also been studied; in particular, an upper bound using the parameter feedback edge set number $f$ and the number $\ell$ of leaves of the graph, was given in [20, 15]. This was improved to $\operatorname{meg}(G)\leq 3f+\ell+1$ (which is tight up to an additive factor of 1) in [7]. The relation between $\operatorname{meg}(G)$ and other various closely related geodetic-parameters is studied in [18].
@@ 2 Polynomial-time algorithm for interval graphs
We begin with some preliminary notation. A graph is an interval graph if we can find a set of intervals on the real line such that each vertex is assigned an interval and two vertices are adjacent if and only if their corresponding intervals intersect. Consider $G$ to be an interval graph. We define $I(a)$ to be the interval representation of a vertex $a$ in $G$. Let $l_{a}$ (resp., $r_{a}$) be the left (resp., right) end point of $I(a)$. Since $G$ is an interval graph, we can consider two total orderings on $G$, namely $<_{l}$ and $<_{r}$ as follows. We say $a<_{l}b$ (resp., $a<_{r}b$) if $l_{a}<l_{b}$ (resp., $r_{a}<r_{b}$) for any two vertices $a,b$ in $G$. Similarly, we denote $a=_{l}b$ (resp., $a=_{r}b$) when $l_{a}=l_{b}$ (resp., $r_{a}=r_{b}$) Let $\ell^{*}$ be a vertex such that $I(\ell^{*})$ is minimal with respect to $<_{l}$. Similarly, let $r^{*}$ be a vertex whose $I(r^{*})$ is maximal with respect to $<_{r}$. Refer to [25] for an algorithm for finding an interval representation of an interval graph in polynomial time.
@@
If a vertex $v$ of a graph $G$ has the property mentioned in Lemma 2.1, then we say that $v$ has property P, and call such a special vertex $u$ (as per Lemma 2.1) as support of $v$. For a vertex subset $S$ of $G$, we define the set,
@@
Let $v$ be a vertex in a graph $G$. If $v$ has property $P$ in $G$, then $d_{G-v}(N(v))\leq 4$. Moreover, if $G$ is an interval graph, then the converse is also true.
@@
Let the vertex $v$ have property $P$ in $G$. Then by Lemma 2.1 the set $N(v)$ contains a support $u$ of $v$. Let $x_{1}$ and $x_{2}$ be two vertices, distinct from $u$, of $N(v)$. Notice that, either $x_{i}$ is adjacent to $u$, or, by Lemma 2.1, has a $2$-path $uw_{i}x_{i}$, for $i\in\{1,2\}$ with $w_{i}\neq v$. In either case, we have $d_{G-v}(u,x_{i})\leq 2$, and thus $d_{G-v}(x_{i},x_{j})\leq 4$.
Next let us assume that $G$ is an interval graph, and consider any interval representation of $G$. We want to show that the converse is also true in this case. Therefore, for some vertex $v$ in $G$, suppose that $d_{G-v}(N(v))\leq 4$. We have to show that $v$ has property $P$.
Notice that, if there exists another vertex $u$ satisfying $I(v)\subseteq I(u)$, then by Lemma 2.1 $v$ has property $P$ where $u$ is a support of $v$. Therefore, we can assume that there is no vertex $u$ in $G$ satisfying $I(v)\subseteq I(u)$. Let $a$ be minimal with respect to $<_{r}$ among all neighbors of $v$ and let $b$ be maximal with respect to $<_{l}$ among all neighbors of $v$. If $v_{r}\leq a_{r}$, then it implies that all intervals corresponding to each neighbor of $v$ intersect the point $a_{r}$, and hence $v$ has property $P$. Thus we may assume that $a_{r}<v_{r}$. Similarly, we may assume that $v_{l}<b_{l}$.
If $b_{l}\leq a_{r}$, then any neighbor $u$ of $v$ will contain the interval $[b_{l},a_{r}]$ due to the definitions of $a$ and $b$. In particular, $a$ and $b$ will dominate $N(v)$, thus implying $v$ has property $P$ where $a$ plays the role of its support. Thus we may assume that $a_{r}<b_{l}$.
@@
If $d_{G-v}(N(v))=1$, then $N(v)$ is a clique. This implies that $v$ has property $P$ with any vertex of $N(v)$ as support.
If $d_{G-v}(N(v))\geq 2$, then due to the choices of $a$ and $b$, for any vertex $x\in N(v)$ we have $I(x)$ intersecting the interval $[a_{r},b_{l}]$. In particular, if $d_{G-v}(N(v))=2$, then $[a_{r},b_{l}]\subseteq I(u_{1})$. Thus, $u_{1}$ is adjacent to all the vertices of $N(v)$, implying $v$ has property $P$ with support $u_{1}$.
If $d_{G-v}(N(v))=3$ or $4$, then we will show that $v$ has property $P$ where $u_{2}$ is the support of $v$. Let $u_{2}vx$ be an induced $2$-path of $G$. Note that,
@@
Thus, $x$ must be adjacent to, at least one of $u_{1}$ and $u_{3}$ (if it exists). Without loss of generality, suppose that $x$ is adjacent to $u_{1}$. Thus, the $2$-path $u_{2}vx$ is part of the $4$-cycle $u_{2}vxu_{1}u_{2}$. Therefore, we can conclude that $v$ has property $P$. ∎
@@
There exists an interval representation of $G$ such that $\ell^{*}$ and $r^{*}$ have property $P$.
@@
Consider any interval representation of $G$. Suppose there is no $x\in V(G)$ such that $I(x)\subseteq I(\ell^{*})$, then we know that $N(\ell^{*})$ is a clique. Hence from [17] and from Lemma 2.1, we can infer that $\ell^{*}$ has property $P$. Suppose on the other hand, that there are some vertices in $G$ whose corresponding interval is entirely contained in the interval of $\ell^{*}$. Let the set $K=\{x\in V(G):I(x)\subseteq I(\ell^{*})\}$. As $K$ is non-empty, let $v$ be the minimal element of $K$ with respect to $<_{l}$. So there does not exist any vertex $t\in N(v)$, such that $\ell^{*}vt$ is an induced $2$-path implying $\ell^{*}$ is a support of $v$. Hence $v$ has property $P$. In this case, we alter the interval representation of $v$ by extending it towards the left such that $v$ becomes the minimal element of $G$ with respect to $<_{l}$. Notice that, this is also an interval representation of the same graph. However, in this interval representation of $G$, $v$ plays the role of $\ell^{*}$.
The proof of $r^{*}$ having property $P$, subject to slight modification of the interval representation of $G$, is similar. ∎
@@
Let $G$ be an interval graph. Then the set of all vertices of $V(G)$ with the property $P$ forms an MEG-set of $G$. Moreover, MEG-set can be solved in polynomial time on interval graphs.
@@
Let $M$ be the set of all vertices in $G$ with the property $P$. We will first prove that $M$ is an (optimal) MEG-set of $G$, and later show how to compute $M$ in polynomial time. That is, if $pq\in E(G)$, we have to prove that $pq$ is monitored by two vertices in $M$. If $p,q\in M$, then the edge $pq$ is trivially monitored by the vertices $p$ and $q$. Thus, without loss of generality let us assume that $q\not\in M$. Also let us fix a particular interval representation of $G$ for which we have $l^{*},r^{*}\in M$, for the rest of the proof. It is possible to choose and fix such an interval representation due to Lemma 2.3.
Observe that, if $I(q)\subseteq I(u)$ for some vertex $u$, then the vertex $u\in N(q)$ is adjacent to all vertices in $N(q)$ and plays the role of a support of $q$. In particular, this would imply that $q$ has property $P$, and thus, $q\in M$, contradicting our assumption. Therefore, there is no vertex in $G$ whose interval contains $I(q)$ as a sub-interval.
@@
Next we show that the vertex $q$ belongs to the path $P$. If not, since the length of the subpath $P_{q}$ is $2$ or less, then $d_{G-q}(N(q))\leq 4$, which is impossible due to Corollary 2.2 since $q$ does not have property $P$, a contradiction. Therefore, we can conclude that $q$ is a vertex of the path $P$.
@@
For the moreover part, note that due to Lemma 2.1 we know $M$ is the unique minimum MEG-set of $G$. Thus, if we can find out which vertices of $G$ belong to $M$ in polynomial time, we will be done. Observe that, using the moreover part of Corollary 2.2, it is possible to decide whether a given vertex of an interval graph $G$ has property $P$ or not in polynomial time, by checking, for every vertex $v$ of $G$, the maximum distance within vertices of $N(v)$ in the subgraph $G-v$. ∎

### v2

@@ 1 Introduction
In the introductory paper on MEG-sets [20] (whose extended version is [15]), the parameter $\operatorname{meg}(G)$ was determined for various graph families such as trees, cycles, unicyclic graphs, complete graphs, complete multipartite graphs, hypercubes and grids. This parameter was further studied for Mycielskian graph classes [30], line graphs [3], Cartesian and strong products of graphs [24], corona products [15], cluster, lexicographic product, direct product and the join operation [35]. The effects of clique sum and subdivision on $\operatorname{meg}(G)$ are also explored in [18]. The relation between $\operatorname{meg}(G)$ and other parameters of graphs has also been studied; in particular, an upper bound using the parameter feedback edge set number $f$ and the number $\ell$ of vertices of degree $1$ of the graph, was given in [20, 15]. This was improved to $\operatorname{meg}(G)\leq 3f+\ell+1$ (which is tight up to an additive factor of 1) in [7]. The relation between $\operatorname{meg}(G)$ and other various closely related geodetic-parameters is studied in [18].
@@ 2 Polynomial-time algorithm for interval graphs
We begin with some preliminary notation. A graph is an interval graph if we can find a set of intervals on the real line such that each vertex is assigned an interval and two vertices are adjacent if and only if their corresponding intervals intersect. Consider $G$ to be an interval graph. We define $I(a)$ to be the interval representation of a vertex $a$ in $G$. Let $l_{a}$ (resp., $r_{a}$) be the left (resp., right) end point of $I(a)$. Since $G$ is an interval graph, we can consider two total orderings on $G$, namely $<_{l}$ and $<_{r}$ as follows. We say $a<_{l}b$ (resp., $a<_{r}b$) if $l_{a}<l_{b}$ (resp., $r_{a}<r_{b}$) for any two vertices $a,b$ in $G$. Similarly, we denote $a=_{l}b$ (resp., $a=_{r}b$) when $l_{a}=l_{b}$ (resp., $r_{a}=r_{b}$). Let $\ell^{*}$ be a vertex such that $I(\ell^{*})$ is minimal with respect to $<_{l}$. Similarly, let $r^{*}$ be a vertex whose $I(r^{*})$ is maximal with respect to $<_{r}$. Refer to [25] for an algorithm for finding an interval representation of an interval graph in polynomial time.
@@
If a vertex $v$ of a graph $G$ has the property mentioned in Lemma 2.1, then we say that $v$ is a mandatory vertex, and call such a “special” vertex $u$ (that is, for any induced $2$-path $uvx$, where $x\in N(v)$, is part of a $4$-cycle) a support of $v$. Observe that, if for no $x\in N(v)$, the path $uvx$ is an induced $2$-path, then $u$ is a support. This is, for instance, the case when $v$ is simplicial. For a vertex subset $S$ of $G$, we define the set,
@@
Let $v$ be a vertex in a graph $G$. If $v$ is a mandatory vertex in $G$, then $d_{G-v}(N(v))\leq 4$. Moreover, if $G$ is an interval graph, then the converse is also true.
@@
Let $v$ be a mandatory vertex in $G$. Then by Lemma 2.1 the set $N(v)$ contains a support $u$ of $v$. If $|N(v)|\leq 2$, then either $v$ is simplicial and we are done, or there exists $x_{1}\in N(v)$ distinct from $u$. By definition of a support, $d_{G-v}(x_{1},u)\leq 2$ and we are done as well. Let $x_{1}$ and $x_{2}$ be two vertices, distinct from $u$, of $N(v)$. Notice that, either $x_{i}$ is adjacent to $u$, or, by Lemma 2.1, has a $2$-path $uw_{i}x_{i}$, for $i\in\{1,2\}$ with $w_{i}\neq v$. In either case, we have $d_{G-v}(u,x_{i})\leq 2$, and thus $d_{G-v}(x_{i},x_{j})\leq 4$.
Next let us assume that $G$ is an interval graph, and consider any interval representation of $G$. We want to show that the converse is also true in this case. Therefore, for some vertex $v$ in $G$, suppose that $d_{G-v}(N(v))\leq 4$. We have to show that $v$ is a mandatory vertex of $G$.
Notice that, if there exists another vertex $u$ satisfying $I(v)\subseteq I(u)$, then by Lemma 2.1 $v$ is a mandatory vertex where $u$ is a support of $v$. Therefore, we can assume that there is no vertex $u$ in $G$ satisfying $I(v)\subseteq I(u)$. Let $a$ be minimal with respect to $<_{r}$ among all neighbors of $v$ and let $b$ be maximal with respect to $<_{l}$ among all neighbors of $v$. If $v_{r}\leq a_{r}$, then it implies that all intervals corresponding to each neighbor of $v$ intersect the point $a_{r}$, and hence $v$ is a mandatory vertex. Thus we may assume that $a_{r}<v_{r}$. Similarly, we may assume that $v_{l}<b_{l}$.
If $b_{l}\leq a_{r}$, then any neighbor $u$ of $v$ will contain the interval $[b_{l},a_{r}]$ due to the definitions of $a$ and $b$. In particular, $a$ and $b$ will dominate $N(v)$, thus implying $v$ is a mandatory vertex where $a$ plays the role of its support. Thus we may assume that $a_{r}<b_{l}$.
@@
If $d_{G-v}(N(v))=1$, then $N(v)$ is a clique. This implies that $v$ is a mandatory vertex with any vertex of $N(v)$ as support.
If $d_{G-v}(N(v))\geq 2$, then due to the choices of $a$ and $b$, for any vertex $x\in N(v)$ we have $I(x)$ intersecting the interval $[a_{r},b_{l}]$. In particular, if $d_{G-v}(N(v))=2$, then $[a_{r},b_{l}]\subseteq I(u_{1})$. Thus, $u_{1}$ is adjacent to all the vertices of $N(v)$, implying $v$ is a mandatory vertex with support $u_{1}$.
If $d_{G-v}(N(v))=3$ or $4$, then we will show that $v$ is a mandatory vertex where $u_{2}$ is the support of $v$. Let $u_{2}vx$ be an induced $2$-path of $G$. Note that,
@@
Thus, $x$ must be adjacent to, at least one of $u_{1}$ and $u_{3}$ (if it exists). Without loss of generality, suppose that $x$ is adjacent to $u_{1}$. Thus, the $2$-path $u_{2}vx$ is part of the $4$-cycle $u_{2}vxu_{1}u_{2}$. Therefore, we can conclude that $v$ is a mandatory vertex. ∎
@@
There exists an interval representation of $G$ such that $\ell^{*}$ and $r^{*}$ are mandatory vertices.
@@
Consider any interval representation of $G$. Suppose there is no $x\in V(G)$ such that $I(x)\subseteq I(\ell^{*})$, then we know that $N(\ell^{*})$ is a clique. Thus, $\ell^{*}$ is a simplicial vertex, hence from Lemma 2.1 of [17] and from Lemma 2.1, we can infer that $\ell^{*}$ is a mandatory vertex. Suppose on the other hand, that there are some vertices in $G$ whose corresponding interval is entirely contained in the interval of $\ell^{*}$. Let the set $K=\{x\in V(G):I(x)\subseteq I(\ell^{*})\}$. As $K$ is non-empty, let $v$ be the minimal element of $K$ with respect to $<_{l}$. So there does not exist any vertex $t\in N(v)$, such that $\ell^{*}vt$ is an induced $2$-path implying $\ell^{*}$ is a support of $v$. Hence $v$ is a mandatory vertex. In this case, we alter the interval representation of $v$ by extending it towards the left such that $v$ becomes the minimal element of $G$ with respect to $<_{l}$. Notice that, this is also an interval representation of the same graph. However, in this interval representation of $G$, $v$ plays the role of $\ell^{*}$.
The proof of $r^{*}$ is a mandatory vertex, subject to slight modification of the interval representation of $G$, is similar. ∎
@@
Let $G$ be an interval graph. Then the set of all mandatory vertices of $V(G)$ forms an MEG-set of $G$. Moreover, MEG-set can be solved in polynomial time on interval graphs.
@@
Let $M$ be the set of all vertices in $G$ that are mandatory vertices. We will first prove that $M$ is an (optimal) MEG-set of $G$, and later show how to compute $M$ in polynomial time. That is, if $pq\in E(G)$, we have to prove that $pq$ is monitored by two vertices in $M$. If $p,q\in M$, then the edge $pq$ is trivially monitored by the vertices $p$ and $q$. Thus, without loss of generality let us assume that $q\not\in M$. Also let us fix a particular interval representation of $G$ for which we have $l^{*},r^{*}\in M$, for the rest of the proof. It is possible to choose and fix such an interval representation due to Lemma 2.3.
Observe that, if $I(q)\subseteq I(u)$ for some vertex $u$, then the vertex $u\in N(q)$ is adjacent to all vertices in $N(q)$ and plays the role of a support of $q$. In particular, this would imply that $q$ is a mandatory vertex, and thus, $q\in M$, contradicting our assumption. Therefore, there is no vertex in $G$ whose interval contains $I(q)$ as a sub-interval.
@@
Next we show that the vertex $q$ belongs to the path $P$. If not, since the length of the subpath $P_{q}$ is $2$ or less, then $d_{G-q}(N(q))\leq 4$, which is impossible due to Corollary 2.2 since $q$ is not a mandatory vertex, a contradiction. Therefore, we can conclude that $q$ is a vertex of the path $P$.
@@
For the moreover part, note that due to Lemma 2.1 we know $M$ is the unique minimum MEG-set of $G$. Thus, if we can find out which vertices of $G$ belong to $M$ in polynomial time, we will be done. Observe that, using the moreover part of Corollary 2.2, it is possible to decide whether a given vertex of an interval graph $G$ is a mandatory vertex or not in polynomial time, by checking, for every vertex $v$ of $G$, the maximum distance within vertices of $N(v)$ in the subgraph $G-v$. ∎